OpenStudy (anonymous):

If h(x) = e^x, then h^-1 (x) = ? A)In(x) or B)log(x) or C)10x or D) 1/e^x

OpenStudy (solomonzelman):

y=e^x x=e^y ln(x)=ln(e^y) ln(x)=y

OpenStudy (anonymous):

So A, B, C, D? Thanks! @SolomonZelman